OpenStudy (anonymous):

Describe the differences between the graph of y= -3/4(x-6)^2+4 and the standard position graph of y = x^2

OpenStudy (anonymous):

alg. 2 btw

OpenStudy (anonymous):

y = x^2: Vertex at (0, 0) Focus: ((0, 1/4) Opens upward y= -3/4(x-6)^2+4: Vertex at (6, 4) Focus: (0, -1/3) Opens downward. Wider than y = x^2

OpenStudy (anonymous):

okay. now how do i describe the first equation that was given?

OpenStudy (anonymous):

I kind of just did. You just have to form sentences using it.

OpenStudy (anonymous):

ohhh lol i see it now. at first i could not see it clear

OpenStudy (anonymous):

may you help me on one more. i tried this on and i cannot seem to get it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 – 12x + 22?

OpenStudy (anonymous):

Use the formula x = -b/2a

OpenStudy (anonymous):

what is -b= to and a=to

OpenStudy (anonymous):

a is 2 and b is -12 Therefore, -(-12)/2(2) = 12/4 = 3 The equation of the line of symmetry is x = 3

OpenStudy (anonymous):

thank you so much !